Session: Mar 2008 (session 1)
Wavelength: 6 cm


To give feedback on this experiment, click here

Experiment Feedback

Urumqi (Wenjun Yang at Tue Mar 11 12:47:13 2008) Success :

Torun (K. Borkowski at Tue Mar 11 13:01:12 2008) Success :
No known problems (apart of phase jumps discovered by Bob earlier, which may be still present).

Medicina (C.Bortolotti at Tue Mar 11 13:09:53 2008) Success :
No knowed problems

Onsala (Michael Lindqvist at Tue Mar 11 15:04:58 2008) Success :
No known problems apart from bad weather (rain).

Westerbork (Tony Foley at Tue Mar 11 15:39:12 2008) Minor failures : PHAS
Wb used in phased array mode. Real time control system died before the end. No useful data after 11:43UT

Effelsberg (Uwe Bach at Wed Mar 12 09:28:34 2008) Success :

Noto (P. Cassaro at Thu Mar 13 12:46:13 2008) Success :

Jodrell_Bank (Alastair Gunn at Wed Mar 19 14:54:43 2008) Minor failures : TSYS
The Lovell Telescope was used (Jb1). Noise diode relay switch was probably in the wrong position, so tpi values are meaningless. No other known problems.

Cambridge (Alastair Gunn at Wed Mar 19 14:55:35 2008) Success :
No known problems.